What is the probability of picking a heart or diamond from a pack of playing cards?
number of cards in a deck =52
number of cards that are heart =13
number of cards that are diamond =13
Probability that the card drawn is a heart or a diamond=13/52 +13/52 =26/52 or 1/2
How many 3 letter arrangements are possible for the word bookkeeper?
Permutations of 10 letters taken 3 a time = 10 x 9 x 8 = 720
What is the theoretical probability of randomly selecting a 3 child family with 3 girls?
The gender of children is not a constant. It depends on the parental predispositions - mainly on the father's genes but apparently the mother's age can also affect the gender. The details of these influences are not understood well enough to develop an accurate model.
If you make the simplistic but unjustified assumption that both genders are equally likely, the answer is 1/8.
When all outcomes have the same probability of occurring the outcomes are?
"Equally likely"; nothing more, nothing less.
What is the probability of getting at least one head on three tosses of a coin?
In the experiment of tossing a fair coin 3 times, the sample space is made of the
following 8 equiprobable events: S = {HHH, THH, HTH, HHT, TTH, THT, HTT, TTT}.
The events that at least have one head are 7. So the probability of getting at least
one head on three tosses is: P(at least one H) = 7/8 = 0.875 = 87.5%
How many permutations are in the word united?
UNITED = 6 letters
The letters in the word UNITED did not repeat so the
number of permutations = 6!
= 6x5x4x3x2
=720
How many even sums can you get from rolling two dice?
You can get 2, 4, 6, 8, 10, and 12, six even sums all together.
What is the probability of picking a jack and a diamond from a deck of cards?
1/52. Only one card, the Jack of diamonds, will satisfy your requirements.
Is it possible that in standard normal distribution standard deviation known and mean unknown?
it is possible to distribute standard deviation and mean but you dont have to understand how the mouse runs up the clock hicorky dickory dock.
How many ways can 9 books be arranged on a shelf?
Assuming that each of the nine books is unique, that there are only nine positions open on the given shelf, and that each book can fit in each position, the answer is 9! (nine factorial) which is equal to 362,880.
If the nine books are not all unique (i.e. there are two of the same book), the number should decrease taking into effect that reversing the positions of the identical books makes no overall change.
If there is one additional position, the number should increase to 10! or 3,628,800. If there is more than one additional position, the number will still increase, but not to 11! or 12! since both open spaces are reversible just like the identical books.
How many ways can you choose 2 books from the 6 on the shelf?
20, without replacing the first book chosen.
25 with replacing it.
How many outcomes are possible if a number cube is rolled twice?
The possible outcomes for rolling a number cube twice are:
1,1 1,2 1,3 1,4 1,5 1,6 2,1 2,2 2,3 2,4 2,5 2,6 3,1 3,2 3,3 3,4 3,5 3,6 4,1 4,2 4,3 4,4 4,5 4,6 5,1 5,2 5,3 5,4 5,5 5,6 6,1 6,2 6,3 6,4 6,5 6,6
This is a total of 36 different combinations.
What is the outcome if two coins are tossed and one number cube is rolled?
Let's call one coin A and the other B. omes
The possible outcomes for the coins are; A heads and B tails, A tails and B heads, A and B heads, A and B tails. That's four outcomes.
The possible outcomes for a single die (as in dice) are six since a die has six faces,
So four times six is twenty four possible outcomes.
What is the probability of obtaining exactly 4 tails from the 9 flips of the coin?
It is approx 0.2461
What happens to theoretical and experimental probability when you increase the number of trials?
When you increase the number of trials of an aleatory experiment, the experimental
probability that is based on the number of trials will approach the theoretical
probability.
Is the data is sample or population?
Data is neither sample nor population. Data are collected for attributes. These can be for a sample or a population.
What is the probability of that you roll a multiple of 3 or 5 on a number cube?
The probability that you roll a multiple of 3 (3 and 6) in a fair die is: P(3 or 6) = 2/6
=1/3 = 0.333... ≈ 33.3%.
The probability that you roll a multiple of 5 (5) is: P(5) = 1/6.
The probability that you roll a multiple of 3 or 5 is: P(3 or 6 or 5) = 2/6 + 1/6 = 1/2 = 0.50 = 50%
What is the probability of not drawing a heart?
Number of cards in a deck = 52
Number of cards that are heart = 13
Therefore number of cards that are not heart = 52-13 = 39
Probability of not drawing a heart = 39/52 or 3/4